OpenStudy (anonymous):

Sam has a total of 58 salmon. If the number of pink salmon is two more than three times the number of silver salmon, how many pink salmon does he have? 42 14 44 12

OpenStudy (anonymous):

let one of the salmons be x first

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I would let x be the silver salmon then we have the equation: 2+3x+x = 58

OpenStudy (anonymous):

so would i now try to find the value of x

OpenStudy (anonymous):

yep but remember I'm letting x be the silver salmon which is not the final answer u want yet

OpenStudy (anonymous):

ok so when i work it out, I get 4x=56 then i 56/4 = 14 is that correct?

OpenStudy (anonymous):

correct! so u now know how to get the final answer?

OpenStudy (anonymous):

ok so now is subtract 14 from the total 58.

OpenStudy (anonymous):

58 - 14 = 44 answer= 44

OpenStudy (anonymous):

so its the third one, C

OpenStudy (anonymous):

u got it!

OpenStudy (anonymous):

thank you so much for your help Maggie. :0

OpenStudy (anonymous):

anytime Michelle:)